### v1

# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

Abstract
Accurate hydrological understanding and water cycle prediction are crucial for addressing scientific and societal challenges associated with the management of water resources, particularly under the dynamic influence of anthropogenic climate change. Existing reviews predominantly concentrate on the development of machine learning (ML) in this field, yet there is a clear distinction between hydrology and ML as separate paradigms. Here, we introduce physics-aware ML as a transformative approach to overcome the perceived barrier and revolutionize both fields. Specifically, we present a comprehensive review of the physics-aware ML methods, building a structured community (PaML) of existing methodologies that integrate prior physical knowledge or physics-based modeling into ML. We systematically analyze these PaML methodologies with respect to four aspects: physical data-guided ML, physics-informed ML, physics-embedded ML, and physics-aware hybrid learning. PaML facilitates ML-aided hypotheses, accelerating insights from big data and fostering scientific discoveries. We first conduct a systematic review of hydrology in PaML, including rainfall-runoff hydrological processes and hydrodynamic processes, and highlight the most promising and challenging directions for different objectives and PaML methods. Finally, a new PaML-based hydrology platform, termed HydroPML, is released as a foundation for hydrological applications. HydroPML enhances the explainability and causality of ML and lays the groundwork for the digital water cycle’s realization. The HydroPML platform is publicly available at https://github.com/HydroPML.
1 Introduction
Many of the scientific and societal challenges to understanding and preparing for environmental change rest upon our ability to understand and predict water cycle changes [1]. Process-based hydrological models play a crucial role in understanding and managing the Earth’s water resources [2] and planning for water security and managing extremes such as floods and droughts [3]. Many hydrological processes can be described as complex physical dynamics over a wide range of spatial and temporal scales, such as hydrodynamic processes and rainfall-runoff processes. Physical approaches have achieved notable success in explaining and predicting the state changes in a hydrological process [4]. However, process-based hydrology continues to present important challenges [1]. (1) A subset of process-based hydrological methods requires significant computational resources and expertise. For example, a global flood solver requires dramatic computational resources based on a traditional finite difference solver [5]. (2) Process-based hydrological models encounter limitations due to existing knowledge gaps. For example, the two-way feedback between humans and water systems is physically agnostic, but it is crucial for the water cycles [6, 7]. A new paradigm is needed to reduce knowledge gaps and explore the unknowns. (3) Process-based hydrology models frequently struggle to rapidly exploit the information in big data. For example, abundant remote sensing observations and hydrological measurements can be utilized for parameter calibration by reducing the differences between model predictions and these big data. However, most traditional process-based hydrological models rely on iterative parameter calibration, which introduces a significant level of uncertainty. Additionally, the iterative process incurs high computational costs [8].
Machine learning (ML) provides efficient alternatives for learning dynamic geophysical phenomena from massive datasets [9]. Recent works have shown that ML can generate realistic short-term hydrology predictions [10] and significantly accelerate the simulation of hydrodynamic processes [11]. Despite the tremendous progress, ML is purely data-driven by nature, which presents many limitations [12, 13]. (1) The nonlinear and chaotic nature of process-based hydrology poses significant challenges for existing data-driven frameworks, which still adhere to the fundamental rules of statistical inference. (2) ML models often struggle with maintaining physical consistency. In the absence of explicit constraints, data-driven models can produce forecasts that violate the physical laws of the hydrological process, leading to implausible outcomes or plausible, yet unexplainable results. (3) Purely data-driven ML models struggle with generalization, as they face challenges in predicting untrained variables and adapting to unseen scenarios with different distributions. This is especially pronounced in process-based systems, where nonlinearity and changing system parameters contribute to distribution shifts. (4) Purely data-driven models lack interpretability and causality, which compromises the reliability of their projections when circumstances change.
Both ML alone and purely physics-based approaches are inadequate for effectively learning complex dynamic processes, such as hydrodynamic and rainfall-runoff processes. Thus, there is an increasing demand to integrate traditional physics-based approaches with ML models, combining the strengths of both approaches. While existing research on physics-aware ML  [12, 14, 15, 16] is extensive, it is not exhaustive or sufficiently comprehensive to encompass a broad range of research domains related to dynamic processes, particularly hydrological processes.
To our knowledge, there is no systematic summary and analysis of physics-aware ML in general. The ever-growing quantity of physics-aware ML methods makes it difficult to find the proper one for a specific problem of dynamic processes. For example, Lawal et al. [14] study and assess physics-informed neural networks (PINNs) from different researchers’ perspectives and categorize the newly proposed PINN methods into extended PINNs, hybrid PINNs, and minimized loss techniques. A review of the three neural network frameworks [17] (i.e., physics-guided neural networks, PINNs, and physics-encoded neural networks) is presented and analyzed. In addition, Goswami et al. [18] present a review of deep neural operator networks and appropriate extensions with physics-informed deep neural operators. An overview of existing physics-guided deep learning (DL) [12] is provided, and existing physics-guided DL approaches are categorized into physics-guided loss function, physics-guided architecture design, hybrid physics-DL models, and invariant and equivariant DL models. In contrast, Willard et al. [19] categorize existing methodologies into physics-guided ML models and hybrid physics-ML frameworks. However, at present, the information about physics-aware ML in these existing reviews is still limited, and these classification systems are also confusing. Beyond summarization and categorization, insightful analysis of the benefits and limitations of existing physics-aware ML methods can help researchers understand the current research state and trends of the scientific ML community. Thus, systematic analysis of physics-aware ML is crucial to the future development of physical dynamics.
There is also no review of physics-aware ML specific to process-based hydrology. Existing reviews are all related to ML in the field of hydrology, or conceptual reviews of physics-aware ML in geoscience, engineering and environmental systems, or Earth system science [20]. For instance, an introductory review of ML for hydrologic sciences [21] is intended for readers new to the field of ML. Sit et al. [22] provide a comprehensive review of DL approaches used in the water industry for generation, prediction, enhancement, and classification tasks. In addition, Reichstein et al. [23] review the development of ML in the geoscientific context, highlighting DL’s potential to overcome many of the limitations of applying ML. The challenging approaches to combining ML with physical modeling are laid out. Furthermore, differentiable modeling [13] that connects varying amounts of prior knowledge to neural networks (NNs) and trains them together is proposed to offer better interpretability, generalizability, and extrapolation capability for advanced geosciences. However, many hydrological processes can be described as dynamic systems of physical equations, such as shallow water equations in hydrodynamic processes, and mass conservation equations in rainfall-runoff processes. Physics-aware ML is important for hydrology in that it integrates physical principles and domain expertise, and enables accurate predictions and extrapolation in complex hydrological systems. Thus, building a review of physics-aware ML in process-based hydrology is urgently needed for physical hydrology.
There is a significant knowledge gap between physics-aware ML and process-based hydrology. Through extensive literature analysis (Fig. 1(a)), physics-aware ML is developing rapidly in the field of ML through different combinations of physics and ML. For example, partial differential equation (PDE) solutions are represented as neural networks by including the square of the PDE residual in the loss function, which was developed in 1998 [24]. In 2019, this approach was refined further and called PINNs [25], initiating a flurry of follow-up work [26]. In 2022, vanilla PINNs began to be used in the field of hydrodynamics (flood) processes [27, 28, 29]. In the ML field, in order to improve the interpretability, causality, and generalizability of ML, and rapidly capture accurate dynamic processes, dynamic physics has also been refined into physical equations, physical properties, and different physical conditions. These fine-grained physical laws are embedded in the frameworks or modules of ML in different ways for different dynamic problems, employing end-to-end training techniques. However, the current integration of physics in hydrology into ML typically follows a simple data-driven or hybrid learning approach [30], which can lead to increased computational burden and potential biases in hydrological models. Thus, further investigation and analysis of physics-aware ML and hydrology in physics-aware ML are imperative to reduce and elucidate the knowledge gap between physics-aware ML and process-based hydrology.
Bearing these concerns in mind, we make an exhaustive and comprehensive review of the physics-aware ML methods. A structured community of existing methodologies that integrates prior physical knowledge or physics-based modeling into ML is built, and called PaML. As shown in Fig. 1(b), PaML approaches are categorized into four groups based on the way physics and ML are combined: (1) physical data-guided ML: a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data; (2) physics-informed ML: a widely used approach is trained from additional information obtained by enforcing the physical constraints (for example, designing loss functions (regularization)); (3) physics-embedded ML: embedding physics in the model frameworks or modules in an end-to-end manner; and (4) physics-aware hybrid learning: directly combining pure physics-based models, such as numerical methods and hydrology models, with ML models. A systematic analysis of each of these different methods is provided. Then, we conduct a systematic review of process-based hydrology in physics-aware ML, including hydrodynamic processes and rainfall-runoff hydrological processes. Next, PaML-based hydrodynamic modeling and PaML-based rainfall-runoff hydrological forecasts are classified by different objectives and PaML methods. We further highlight the most promising and challenging directions for different application categories, aiming to identify opportunities for advancement and address existing hurdles in the field. Finally, in order to reduce the knowledge gap between physics-aware ML and process-based hydrology, a new PaML-based hydrology platform, termed HydroPML, is proposed as a foundation for applications based on hydrological processes. For example, we apply HydroPML to real-time flood forecasts.
Figure 1: (a) Examine the knowledge gap between physics-aware ML and hydrology in physics-aware ML from the keyword viewpoint. (b) Conceptual framework of physics-aware ML (PaML). PaML includes physical data-guided ML, physics-informed ML, physics-embedded ML, and physics-aware hybrid learning.
The main contributions of this work are summarized below.
  • We conduct an extensive and comprehensive review of physics-aware ML methods, which serve as a transformative approach to bridge process-based hydrology and ML, ultimately facilitating a paradigm shift in both fields. These methods are summarized and categorized into physical data-guided ML, physics-informed ML, physics-embedded ML, and physics-aware hybrid learning.
  • Systematic analyses of four aspects of these PaML methodologies are undertaken, to provide insights and ideas for research within the PaML community.
  • We systematically analyze PaML-based hydrological processes, specifically focusing on hydrodynamic processes and rainfall-runoff hydrological processes. We categorize PaML-based hydrodynamic modeling and PaML-based rainfall-runoff hydrological forecasts based on different objectives and PaML methods. Additionally, we identify the most promising directions and challenges in various application categories, aiming to advance the field and overcome existing obstacles.
  • We release an open platform, termed HydroPML, to reduce the knowledge gap between physics-aware ML and process-based hydrology. In addition, we utilize HydroPML in real-time flood forecasting applications, significantly enhancing water security and fostering resilient water management.
The rest of this paper is organized as follows. Section 2 reviews state-of-the-art physics-aware ML. Section 3 presents process-based hydrology in physics-aware ML (hydroPML), and its applications in real-time flood forecasts.
2 State-of-the-art Physics-aware Machine Learning
Figure 2: The proposed physics-aware machine learning (PaML) community.
Physics-aware ML (PaML) aims to take the best from both physics-based modeling and state-of-the-art ML models to better solve scientific problems. As shown in Fig. 2, a structured community of existing PaML methodologies that integrate prior physical knowledge or physics-based modeling into ML is built. We categorize PaML approaches into four groups based on the way physics and ML are combined, including physical data-guided ML (PDgML), physics-informed ML (PiML), physics-embedded ML (PeML), and physics-aware hybrid learning (PaHL). Each is discussed in detail below.
2.1 Physical Data-guided Machine Learning
PDgML is a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data [54]. PDgML consists of one or a combination of deep neural networks (DNN) [31], convolutional neural networks (CNNs) [32], recurrent neural networks (RNNs) [33], generative adversarial networks (GAN) [34], graph neural networks (GNN) [35], deep reinforcement learning (DRL) [36], Transfomer [37], deep operator networks [38], and physics-discovery neural networks [39]. The objectives and limitations of different physical data-guided ML are briefly discussed in Table 1.
2.1.1 Classical Physical Data-guided Neural Networks
Spatio-temporal modeling. DL has proven to be one of the most powerful machine-learning techniques for solving complex problems using big data. Most of the data generated through smart devices today are time series, and their prediction is one of the most frequent and current problems in almost all research areas [40]. The Elman network, the first RNN, incorporated the $t$ state of a hidden unit to make predictions in data sequences [41]. Long short-term memory (LSTM) [42] recurrent networks emerge to solve the vanishing gradient problem in RNNs, which uses three gates to keep long-standing relevant information and discard irrelevant information. Physics-guided LSTM [43] is designed for the extreme events and monotonic relationships of a deep learning network for rainfall-runoff simulation. Three physical mechanisms are therefore considered to constrain the LSTM network from high flow, low flow, and monotonic properties. However, existing methods are designed under limited problem settings, like predicting 48 points or fewer [44]. Recent studies have shown the potential of Transformer to increase prediction capacity. Thus, Informer [45] is proposed to enhance the prediction capacity in the long sequence time-series forecasting problem, which validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequence time-series outputs and inputs.
Additionally, an army of physical data-guided spatio-temporal models [46, 47, 48] is proposed by imposing temporal and spatial inductive biases based on RNNs and CNNs. For example, ConvLSTM [49] is a type of recurrent neural network for spatio-temporal prediction that has convolutional structures in both the input-to-state and state-to-state transitions. Furthermore, inspired by the strong representation capacity of Transformer architectures (such as bidirectional encoder representations from transformers (BERT) [50] in natural language processing, vision Transformer in computer vision [51, 52]), quite a few physics-guided Transformer-based frameworks are presented. For example, a space-time Transformer for Earth system forecasting, Earthformer [53], is presented based on an efficient space-time attention block. A global data-driven weather forecasting model, FourCastNet [54], is proposed to provide accurate short- to medium-range global predictions at 0.25°$\degree$ resolution. In addition, in order to improve prediction accuracy and uncertainty, irregular space-time grid prediction problems, and other issues, a collection of classic spatio-temporal models are proposed, such as GraphCast [55] based on GNN for medium-range global weather forecasting, physically constrained GANs for improving local distributions and spatial structure of precipitation fields [56], and spatially-irregular forecasting based on geometric deep learning [57].
Dynamic system modeling. Simulating multi-dimensional PDE systems with data-driven NNs has been a renewed research topic, that dates back to the last century [58]. Physical data-guided machine learning for dynamic system modeling is designed to learn a PDE model (if one exists) from a given data set and perform accurate and efficient predictions using the learned model. Development has taken place in roughly three phases.
First, any early attempt [59] at the data-driven discovery of hidden physical laws compares numerical differentiations of the experimental data with analytic derivatives of candidate functions and applies symbolic regression and an evolutionary algorithm to determine the nonlinear dynamic system. Recently, motivated by the latest development of neural networks, a fully non-mechanistic method, PDE-Net [60], is designed to accurately predict the dynamics of complex systems and uncover the underlying hidden PDE models based on learning convolution kernels (filters). Further works have extended the approach with residual CNNs [61], symbolic neural networks [62], feed-forward networks [63], and deep fully convolutional networks [64]. However, these models are fundamentally limited to discretizing the input domain with a sample-inefficient grid, rendering them unable to handle temporally or spatially sparse or non-uniform observations commonly encountered in realistic applications. Recently, some convolution operators, such as Mesh-Conv [65], have been proposed to model non-uniform structured or unstructured spatial mesh.
Second, neural message passing-based models such as GNN-based model [66], autoregressive solver [67], and contrastive learning-based message passing graph neural networks (MP-GNNs) [68] are related to interaction networks, where the object’s state evolves as a function of its neighboring objects, which forms dynamic relational graphs instead of grids. Neural message passing is a general framework for supervised (physical data-guided) learning on graphs [69]. In addition to the common mesh-based neural message passing-based approaches, particle-based frameworks have been attracting a lot of interest in scientific modeling for physical systems, especially those employing GNNs [70, 71, 72]. Although neural message passing-based models can achieve geometry-adaptative learning of nonlinear PDEs with arbitrary domains, these models can apply message-passing between small-scale moving and interacting objects.
Third, DRL has recently been adopted in a wide range of data-driven dynamic system modeling for its ability to solve decision-making problems that were previously out of reach due to a combination of non-linearity and high dimensionality [73, 74]. For example, a DRL-based method [75] is proposed to control a coupled 2D system involving both fluid and rigid bodies, making use of a position-based reward function.
2.1.2 Deep Operator Networks
Operator learning strategies are predicated on the principles of the universal approximation theorem [76]. This approximation ability has formed the basis of modern deep operator network development. Here, we present an introduction to deep operator networks (DeepONet) [76], the Fourier neural operator (FNOs) [77], the wavelet neural operators (WNOs) [78], and the graph neural operators (GNOs) [79].
DeepONets utilize the aforementioned branch-trunk architecture in order to map finite inputs to the response space [76]. This enables domain encoding to be handled by the trunk network, with any number of branch networks used to handle the various inputs associated with the problem. For example, in [80], knowledge of the dynamics of decaying evolution of the Brusselator reaction-diffusion system is incorporated by employing a trigonometric feature expansion of the temporal input of the trunk network. In addition, to employ DeepONet for realistic setups (multiple input functions) and diverse applications, a multiple input DeepONet [81] is presented for multiple initial conditions and boundary conditions at the same time. In addition, to enable the learning of the DeepONet with a smaller set of parameters, HyperDeepONet [82] is proposed by using the expressive power of the hyper network. However, several studies point out that the slow decay rate of the lower bound leads to inaccurate approximation operator learning for complex physical dynamics using DeepONet [83, 84].
FNOs [77, 85] use Fourier Transformations to move to the infinite-dimensional response space. The operator takes input functions specified on a well-defined equally spaced lattice grid and outputs the field of interest on the same grid points. The network parameters are defined and learned in the Fourier space rather than in the physical space. FNO has become one of the mainstream physical data-guided deep operator networks used to solve PDEs [38]. Specifically, to solve PDEs on arbitrary geometries, Geo-FNO [86] is designed by learning to deform the input (physical) domain. Domain agnostic FNO [87] is proposed by incorporating a smoothed characteristic function in the integral layer architecture of FNOs. However, the FNO and the geo-FNO perform worse and are unstable on complex geometries and noisy data. Thus, factorized FNO [88] is proposed to improve representation stability for the operator. In addition, a host of variants of FNOs are proposed to improve the generalization of FNOs for solving PDEs, such as incremental FNO [89].
One major shortcoming of FNOs is that the basis functions in FFTs are generally frequency localized with no spatial resolution [90]. Thus, WNOs are proposed, to learn the network parameters in the wavelet space that are both frequency and spatial localized, and thereby learn the signal patterns more effectively. Recently, a coupled multiwavelets neural operator learning scheme [91] is proposed to solve coupled PDEs, by decoupling the coupled integral kernels in the wavelet space. This approach has shown that WNO can handle domains with both smooth and complex geometries, and has been applied in learning solution operators for a highly nonlinear family of PDEs with discontinuities and abrupt changes in the solution domain and boundary.
GNOs [79] are computed by message passing on graph networks, which are able to capture the graph (non-local) structure of the data. Graph kernel networks (GKNs) are mainly used in GNOs. The idea of GKNs comes from parameterizing Green’s function in an iterative architecture [79]. As an integral neural network similar to FNOs, GKNs are composed of a lifting layer, iterative kernel integration layers, and a projection layer. However, it has been found that GKNs might be unstable when increasing the number of kernel integration layers [92]. Thus, a nonlocal neural operator [92] is proposed by using a resolution-independent nonlocal kernel network.
2.1.3 Physics-discovery Neural Networks
The discovery of dynamic behavior is crucial, as it enables us to comprehend the underlying physical laws that regulate complex systems. However, discovering governing equations is a challenging task due to complex and nonlinear systems, as well as noisy and incomplete data. With the rapid advance of machine intelligence [12], the past three decades have seen a considerable amount of physics-discovery neural networks growing up around the interplay between ML and scientific discovery. Essentially, two genres have been observed based on unavailable and available prior knowledge of the physics system. First, the natural and optimal solution to automatically identifying the governing equations for dynamic systems is to learn a symbolic model from experimental data. Hence, symbolic regression [93] and symbolic neural networks (SNNs) [94, 95, 96] have been explored for inferring concise equations without any prior knowledge. A physics-inspired method called AI Feynman [97, 98] is proposed for symbolic regression. AI Feynman employs NNs to discover generalized symmetries in complex data, enabling the recursive decomposition of difficult problems into simpler ones with fewer variables. GNN [99] is also utilized to identify the nontrivial relation due to its excellent inductive capability. In addition, PDgML can apply for knowledge discovery that cannot be modeled from a physical process-based perspective, such as the two-way feedback between human and water systems [6, 7]. Second, discovering dynamic behavior from data is often accomplished by defining a large set of possible mathematical basis functions or process-based models based on prior knowledge. The most representative works are SINDy [100] and PDE-FIND [101] for ordinary differential equations (ODEs) and PDEs, respectively. However, the standard sparse representation-based methods are limited to high-fidelity noiseless measurement, which is usually difficult and expensive to obtain. Hence, many recent efforts have been devoted to discovering PDEs from sparse/noisy data [102, 103].
Table 1: Proposed Physical Data-guided Machine Learning
Categories Types Examples
Objective(s)
Limitation(s)
Classical Physical Data-guided Neural Networks Spatio-temporal modeling LSTM [42]
Three gates to keep long-standing relevant information and discard irrelevant information
Limited generalization, high computational cost, limited physical interpretability
Physics-guided LSTM [43]
A deep learning network for rainfall-runoff simulation
Informer [45]
Long sequence time-series forecasting
ConvLSTM [49]
A spatio-temporal sequence forecasting problem
Earthformer [53]
A space-time Transformer for Earth system forecasting
Geometric deep learning [57]
A spatially-irregular forecast method
Dynamical system modeling PDE-Net [60]
A feed-forward deep network-based PDE solver
Unable to handle temporally or spatially sparse or non-uniform observations
Mesh-Conv [65]
A solver for modeling non-uniform structured or unstructured spatial mesh
High computational cost
Neural message passing-based models [66, 67, 68, 69, 70, 71, 72]
A general framework for supervised learning on graphs (arbitrary domains)
Mesh-based or particle-based; message-passing between small-scale moving and interacting objects
DRL-based method [75]
A coupled 2D control system
Limited capabilities and robustness; unknown behavior and convergence speed in action spaces of high dimensionalities
Deep Operator Networks Deep Operator Networks (DeepONet) [76] Multiple input DeepONet [81]
Employ DeepONet for realistic setups (multiple initial conditions and boundary conditions)
Inaccurate approximation for complex physical dynamics
HyperDeepONet [82]
Enable the learning of the DeepONet with a smaller set of parameters
Fourier neural operator (FNOs) [77] Geo-FNO [86], Domain agnostic FNO [87]
Solve PDEs on arbitrary geometries
Unstable, limited generalization, the basis functions in FFTs are generally frequency localised with no spatial resolution
Factorized FNO [88]
Improve representation stability for the operator
Wavelet neural operatos (WNOs) [78] WNO [78]
Learn the network parameters in the wavelet space
Complexity, limited interpretability
Coupled multiwavelets neural operator [91]
Learn coupled PDEs in the wavelet space
Graph neural operators (GNOs) [79] Graph kernel networks [79]
An integral neural network that parameterizes Green’s function
Difficult to train, computational cost, limited generalization
Nonlocal neural operator [92]
A resolution-independent nonlocal kernel network
Physics-discovery Neural Networks Physics-discovery with unavailable prior knowledge Symbolic neural networks (SNNs) [94, 95, 96]
Infer concise equations without any prior knowledge
Challenging task due to complex and nonlinear systems, as well as noisy and incomplete data
AI Feynman [97, 98]
Discover generalized symmetries in complex data
GNN [99]
Identify the nontrivial relation due to its excellent inductive capability
Knowledge-agnostic discovery [6, 7]
Discover the two-way feedback between human and water systems
Physics-discovery with available prior knowledge SINDy [100]
Discover governing equations from data by sparse identification of nonlinear dynamical systems
Limited high-fidelity noiseless measurement
PDE-FIND [101]
Data-driven discovery for PDEs
2.2 Physics-informed Machine Learning
In scientific computing, physical phenomena are often described using a strong mathematical form consisting of governing differential equations as well as initial and boundary conditions. PiML is a widely-used approach that integrates physical laws into ML models (for example, designing loss functions or regularization), facilitates the accurate capture of dynamic patterns and concomitantly diminishes the search space for model parameters. This approach is sometimes referred to as imposing differentiable constraints in loss functions. It integrates (noisy) data and mathematical models, and implements them through neural networks (physics-informed neural networks) or kernel-based neural operators (physics-informed neural operators). The objectives and limitations of different physics-informed machine learning are briefly discussed in Table 2.2.
Table 2: Proposed Physics-informed Machine Learning
Categories
Types
Examples
Objective(s)
Limitation(s)
Vanilla PINNs
PINNs [25]
A deep learning framework for solving forward and inverse problems involving nonlinear partial differential equations
The challenging optimization landscape from soft physics or PDE constraints; the difficulty of propagating information from the initial or boundary conditions to unseen parts of the interior or to future times; the sensitivity to hyperparameter selection; the slow training/re-optimization process in practical applications and different instances; failing to converge on high-frequency, multi-scale PDEs, or time-dependent PDEs
Problem-specific insights
Universal PINNs [104]
Solve PDEs with a point source by a lower bound constrained uncertainty weighting algorithm and a multi-scale deep neural network
Limited accuracy and generalization
cPINN [105]
Develop a unique conservative physics-informed neural network on discrete domains for nonlinear conservation laws
Limited representation and parallelization capacity
nPINNs [106]
Extend PINNs to parameter and function inference for integral equations such as nonlocal Poisson and nonlocal turbulence models
Challenging optimization; require more residual points
Improved PINNs [107]
An improved PINN approach for localized wave solutions of the derivative nonlinear Schrodinger equation in complex space with faster convergence and optimum simulation performance
Lack of consideration of complex integrable equations
AdaptivePINNs [108]
An automatic numerical solver for the Allen-Cahn and Cahn-Hilliard equations that proposes an improved PINN
Limited application range (focused on the problem of solving differential equations)
CINN [109]
A simple and efficient characteristic-informed neural networks for solving forward and inverse problems involving hyperbolic PDEs
Limited application; limited expression and generalization ability
NSFnets [110]
Physics-informed neural networks for the incompressible Navier-Stokes equations by automatic differentiation
Limited expressivity and generalization; hyperparameters; failure to converge on high-frequency PDEs; high computational cost
GAPINN [111]
Geometry aware physics-informed neural network for solving Navier–Stokes equation on irregular geometries
Limited expressivity of framework; cannot deal with time varying geometries
CPINNs [112]
Coupled physics-informed neural networks for the closed-loop geothermal system
Limited expressivity of framework; hyperparameters
Input data
Bayesian-PINNs [113]
Solve both forward and inverse nonlinear problems described by PDEs and noisy data
Limited representation; lack of large scale dataset testing
ModalPINNs [114]
An extension of physics-informed neural networks with enforced truncated Fourier decomposition for periodic flow reconstruction using a limited number of imperfect sensors
Limited application range
$\Delta$-PINNs [115]
A positional encoding mechanism for PINNs based on the eigenfunctions of the Laplace-Beltrami operator, presented on complex geometries
Limited accuracy using finite elements; computing eigenfunctions of the Laplace-Beltrami operator can become expensive for large meshes
Frameworks
PhyCRNet [116]
An encoder-decoder, ConvLSTM, is proposed for low-dimensional spatial feature extraction and temporal evolution learning. The loss function is defined as the aggregated discretized PDE residuals
Limited accuracy due to finite difference
PINNs with ensemble models [117]
Train an ensemble of PINNs and use the ensemble agreement (confidence) as the criterion for expanding the solution interval to new areas
High computational cost
Physics-informed GNNs [118]
Physics-informed graph neural networks not only provide a representation that is specially tailored to the symmetries and properties of the systems we consider, but they also are naturally invariant to the extent of the system
High computational cost; limited to homogeneous systems
MH-PINNs [119]
Multi-head physics-informed neural network constructed as a potent tool for multi-task learning, generative modeling, and few-shot learning for diverse problems
Limited expressivity of framework; hyperparameters; difficult to optimize
Phase2vec [120]
An embedding method that learns high-quality, physically-meaningful representations of 2D dynamical systems by a convolutional backbone
Failure in the n-dimensional case; dynamics are qualitatively invariant to translations of the input coordinate system
PI-PINN [121]
Orthogonal polynomials are used to construct the neural network for solving nonlinear partial differential equations
Limited expressivity of framework
PhyGeoNet [122]
Physics-constrained CNN learning architecture to learn solutions of parametric PDEs on irregular domains without any labeled data.
Limited accuracy due to finite difference; cannot solve dynamic geometrically irregular domains
Spline-PINN [123]
Approaching PDEs training without any precomputed training data using a physics-informed loss function only; provides fast, continuous solutions that generalize to unseen domains
Limited accuracy due to finite difference
Physics-constrained Unsupervised Learning [124]
Represent meshes naturally as graphs, process them using graph networks, and formulate physics-based loss to provide an unsupervised learning framework for PDEs
Limited accuracy due to finite difference
Improving speed and efficiency
Meta-learning based PINNs
de Avila Belbute-Peres et al. [125] meta-trains a hyper-network that for each task can generate weights for a small neural network. Psaros et al. [126] meta-learns a loss function that is used to optimize the NN. Penwarden et al. [127] proposes a meta-learning approach that learns an initialization of weights such that the NN can be optimized quickly.
Limited expressivity of framework, unsteady in solving related PDEs
XPINNs [128]
A generalized space-time domain decomposition framework, proposed for PINNs to solve nonlinear PDEs on arbitrary complex-geometry domains
Hyperparameters; difficult to optimize
canPINN [129]
Coupled-automatic-numerical differentiation framework that unifies the advantages of automatic differentiation and numerical differentiation, providing more robust and efficient training than automatic differentiation-based PINNs
Limited application scenarios (generally applicable to non-uniform samplings), limited computational efficiency
DT-PINNs [130]
DT-PINNs are trained by replacing these exact spatial derivatives with high-order accurate numerical discretizations computed using meshless radial basis function-finite differences and applied via sparse-matrix vector multiplication
Limited accuracy due to radial basis function-finite differences
Improving generalization
TPINN [131]
Transfer physics-informed neural network for solving forward and inverse problems in nonlinear partial differential equations by parameter sharing
Lack of consideration noise in input data
Multiscale and multiphysics problems
FBPINNs [132]
Finite basis physics-informed neural networks use ideas from domain decomposition to accelerate the learning process of PINNs and improve their accuracy
Fail to coverge on high-frequency PDEs; limited accuracy and expressivity
Curriculum and Sequence-to-sequence PINN [133]
Curriculum regularization and sequence-to-sequence learning, used to address the problems of PINNs that fail to learn relevant physical phenomena for even slightly more complex problems
Limited expressivity of framework; hyperparameters; lack of realistic complex application
Multi-scale PINNs [134]
Spatio-temporal and multi-scale random Fourier features are used to address high-frequency or multi-scale problems in PINN
Hyperparameters
Hyperparameter selection
Self-adaptation PINNs [135]
Self-adaptation PINNs that make weights increase as the corresponding losses increase, which is accomplished by training the network to simultaneously minimize losses and maximize weights
Limited expressivity of framework; the squared residual penalty method imposes a fundamental limitation associated with conditioning
Heuristic method [136]
A heuristic method for optimizing weighted loss functions for solving PDEs with NNs
Data-physics-driven parameter discovery
PINN-TFC [137]
Data-physics-driven parameters discovery of problems modeled via ordinary differential equations
Cannot be applied to the data-driven discovery of problems
PINN-SR
Physics-informed neural network with sparse regression to discover governing partial differential equations from scarce and noisy data for nonlinear spatio-temporal systems
High computational cost need a vast number of collocation points for complex systems; limitations of PINNs
Physics-informed Neural Networks
Discovering governing equations [138]
Solve the inverse problem of parameter identification in discrete, high-dimensional systems inspired by physical applications
Limitations of PINNs; cannot address higher dimensional settings
DeepONet-based
Physics-informed DeepONet [139]
Train the DeepONet by directly incorporating known differential equations into the loss function, along with some labeled datasets
Limitations on fixed grid or basis; cannot achieve discretization invariant; limitation of linear approximation; limited computational efficiency and generalization
FNO-based
PINO [140]
An integration of Fourier operator learning and physics-informed settings
Limited accuracy on solving non-periodic problems; limited adaptability due to the selection of model parameters; it cannot be applied to large-scale and long-term sequence PDE solution
FC-PINO [141]
An architecture that leverages Fourier continuation is proposed to apply the exact gradient method to PINO for nonperiodic problems
Limited adaptability due to the selection of model parameters; it cannot be applied to large-scale and long-term sequence PDE solution; it cannot achieve geometry-adaptive learning
GeoPINS [142]
A geometry-adaptive physics-informed neural solver based on the advantages of no training data in physics-informed neural networks, as well as possessing a fast, accurate, geometry-adaptive and resolution-invariant architecture
Limited adaptability due to the selection of model parameters
Physics- informed Neural Operators
WNO-based
Physics-informed WNO [143]
A physics-informed WNO for learning the solution operators of families of parametric PDEs without labeled training data, based on the advantage of time-frequency localization of wavelets
Complexity; the absence of a standardized procedure to obtain the optimum and desired network architecture of a physics-informed WNO and to assign the weights of boundary/initial condition loss; the lack of information regarding how to learn the operator for systems involving multi-scale physics systems
2.2.1 Physics-informed Neural Networks
PDE solutions are represented as neural networks by including the square of the PDE residual in the loss function, resulting in a neural network-based PDE solver [24]. Recently, this approach is refined further and called “physics-informed neural networks (PINNs) [25],” initiating a flurry of follow-up work. For example, we consider a parametrized PDE system given by
$\begin{gathered}f\left(\mathbf{x},t,u,\frac{\partial u}{\partial\mathbf{x}},\frac{\partial^{2}u}{\partial\mathbf{x^{2}}},\frac{\partial u}{\partial t},\ldots,\vartheta\right)=0,\quad\mathbf{x}\in\Omega,t\in[0,T]\\ u\left(\mathbf{x},t_{0}\right)=g_{0}(\mathbf{x})\quad\mathbf{x}\in\Omega\quad u(\mathbf{x},t)=g_{\Gamma}(t),\quad\mathbf{x}\in\partial\Omega,t\in[0,T],\end{gathered}$ (1)
where $u:[0,T]\times\mathbb{X}\rightarrow\mathbb{R}^{n}$ is the solution, with initial condition $g_{0}(\mathbf{x})$ and boundary condition $g_{\Gamma}(t)$ (which can be periodic, Dirichlet’s, Neumann’s, or other boundary condition). The PDE parameters are denoted by $\vartheta$ = [$\vartheta 1$, $\vartheta 2$, . . . ]. $f$ denotes the residual of the PDE, containing the differential operators (i.e., $\frac{\partial u}{\partial\mathbf{x}},\frac{\partial^{2}u}{\partial\mathbf{x^{2}}},\frac{\partial u}{\partial t}$, . . . ); and $\Omega$ and $\partial\Omega$ represent the physical domain and the boundary, respectively.
Vanilla PINNs [25] directly approximate the solution of differential equations in a physics-informed fashion with less training data [144] or without any training data [145, 110, 28]. In PINNs, solving a PDE system is converted into an optimization problem by minimizing the loss function to iteratively update the NN,
$\begin{gathered}\min_{\theta}\mathcal{L}_{\mathrm{PINN}}(\theta)=w_{i}\mathcal{L}_{IC}+w_{b}\mathcal{L}_{BC}+w_{d}\mathcal{L}_{\text{Data }}+w_{p}\mathcal{L}_{PDE}\\ \mathcal{L}_{IC}=\|u\left(\mathbf{x},t_{0}\right)-g_{0}(\mathbf{x})\|_{\Omega}\\ \mathcal{L}_{BC}=\|u(\mathbf{x},t)-g_{\Gamma}(t)\|_{\partial\Omega}\\ \mathcal{L}_{\text{Data }}=u|_{\Omega}-\hat{u}|_{Data}\\ \mathcal{L}_{PDE}=\|f\left(\mathbf{x},t,u,\frac{\partial u}{\partial\mathbf{x}},\frac{\partial^{2}u}{\partial\mathbf{x^{2}}},\frac{\partial u}{\partial t},\ldots,\vartheta\right)\|_{\Omega},\end{gathered}$ (2)
where $w_{i},w_{b},w_{d},w_{p}$ are regularization parameters that control the emphasis on residuals of initial conditions ($\mathcal{L}_{IC}$), boundary conditions ($\mathcal{L}_{BC}$), training data ($\mathcal{L}_{\text{Data }}$), and PDE ($\mathcal{L}_{PDE}$), respectively. $\mathcal{L}_{\text{data }}$ measures the mismatch between the NN prediction $u$ and the training data $\hat{u}$. The NN parameters are denoted by $\theta$, which takes the spatial and temporal coordinates (x, t), and possibly other quantities, as inputs and then outputs $u$.
Figure 3: (a) Challenges of vanilla PINNs. (b) An example of vanilla PINNs failing to converge on high-frequency and multi-scale PDEs. 1-D convection equations with high-frequency (first row): Train and infer on a spatial-temporary resolution $256\times 100$; 2-D steady incompressible Navier-Stokes equations (second row): Train and infer on a spatial resolution $49\times 77$. Analytical solution and computational fluid dynamics (CFD) represent ground truth.
As shown in Fig. 3(a), in current NN-based ML methods, vanilla PINNs are mainly defined in a pointwise manner. Vanilla PINNs use a neural network as the ansatz of the solution function and optimize a loss function to minimize violation of the given equation by taking advantage of auto-differentiation to compute the exact, mesh-free derivatives. Vanilla PINNs overcome the need to choose a discretization grid that numerical solvers require in many scenarios, such as inverse and higher dimensional problems. However, PINNs are far less efficient than classical methods for solving most PDEs, for many reasons [146, 147, 148, 149]: (1) the challenging optimization landscape from soft physics or PDE constraints; (2) the difficulty of propagating information from the initial or boundary conditions to unseen parts of the interior or to future times; (3) the sensitivity to hyperparameter selection; (4) the slow training/re-optimization process in practical applications and different instances; and (5) failure to converge on time-dependent PDEs or high-frequency, multi-scale PDEs, as shown in Fig. 3(b).
As shown in Table 2.2, a vast and growing body of work aims to address these challenges. First, many problem-specific insights are utilized by improving the training of PINNs. For example, Universal PINNs [104] improve the PINNs by a lower bound constrained uncertainty weighting algorithm and a multi-scale deep neural network. A unique conservative physics-informed neural network (cPINN) [105] is developed on discrete domains for nonlinear conservation laws. Another approach, nPINNs [106] extends PINNs to parameter and function inference for integral equations such as nonlocal Poisson and nonlocal turbulence models. An improved PINN approach [107] is presented for localized wave solutions of the derivative nonlinear Schrodinger equation in complex space, yielding faster convergence and optimum simulation performance. An automatic numerical solver is proposed [108] for the Allen-Cahn and Cahn-Hilliard equations by improved PINNs. A simple and efficient characteristic-informed neural network [109] is developed for solving forward and inverse problems involving hyperbolic PDEs. NSFnets [110] is proposed for the incompressible Navier-Stokes equations by employing automatic differentiation. Furthermore, for solving Navier-Stokes equations on irregular geometries, a geometry-aware physics-informed neural network (GAPINN) [111] is presented. In addition, a coupled physics-informed neural networks (CPINNs) [112] is built for closed-loop geothermal systems.
Second, PINNs are not equipped with input data preprocessing, which may restrict their application, especially for scenarios where the input data are noisy, imperfect, or sparse, or geometry is complex. Bearing these concerns in mind, Bayesian-PINNs [113] make use of both physical laws and scattered noisy measurements to provide predictions and quantify the aleatoric uncertainty arising from the noisy data in the Bayesian framework. An extension of PINNs with enforced truncated Fourier decomposition (ModalPINNs) [114] is proposed for flow reconstruction using imperfect and sparse information. In addition, a positional encoding mechanism for PINNs based on the eigenfunctions of the Laplace-Beltrami operator ($\Delta$-PINNs) [115] is presented on complex geometries.
Third, due to the limited expressive capacity inherent in fully-connected networks within vanilla PINNs, the training process of vanilla PINNs may not be robust and sufficiently stable, and thus convergence to the global minimum cannot be guaranteed. In addition, fully-connected networks struggle to learn multiscale and multiphysics problems. To resolve this issue, more robust NN architectures, such as a multi-scale deep neural network in universal PINNs [104], an ensemble model of PINNs [117], graph neural networks in physics-informed GNNs [118], multi-head physics-informed neural networks (MH-PINNs) [119], a convolutional backbone in Phase2vec [120], and orthogonal polynomials in PI-PINN [121] have been developed. In addition, researchers started to blend physics with RNN-based, CNN-based, or GNN-based learning. For example, a physics-informed convolutional-recurrent learning architecture (PhyCRNet) [116] is proposed for solving PDEs without any labeled data. Physics-constrained CNN learning architecture (PhyGeoNet) [122] is developed to learn solutions of parametric PDEs on irregular domains without any labeled data. A fast and continuous CNN-based solution (Spline-PINN) [123] is designed for approaching PDE training without any precomputed training data. Graph network [124] is used to represent meshes, and physics-based loss is formulated to provide an unsupervised learning framework for PDEs. However, these methods usually have limited accuracy due to gradient solving at the grid scale (such as finite difference).
Fourth, vanilla PINNs tend to be dramatically slower than classic numerical methods. A popular method for rapidly solving problems by PINNs from a distribution of related PDEs is meta-learning based methods. Specifically, de Avila Belbute-Peres et al. [125] meta-train a hypernetwork that for each task can generate weights for a small neural network. The small neural network then becomes the approximate solution to the PDE. Psaros et al. [126] meta-learn a loss function that is used to optimize the NN; this is found to achieve performance benefits in comparison to both hand-crafted and online adaptive loss functions. Penwarden et al. [127] propose a meta-learning approach that, like our Meta-PDE approach, learns an initialization of weights such that the NN can be optimized quickly. Penwarden et al. [127] find that model-agnostic meta-learning (MAML) [150] achieves poor performance, only marginally better than random initialization; the recently published Qin et al. [151] come to the opposite conclusion, that MAML-based Meta-PDE performs well for a given runtime. In addition, a cohort of other methods is also proposed to improve the efficiency of PINNs. For instance, XPINNs [128] is presented by space and time parallelization. To replace the computationally expensive automatic differentiation in PINNs, meshless radial basis function-finite differences in DT-PINNs [130] are applied by sparse-matrix vector multiplication, and a coupled-automatic-numerical differentiation framework (canPINN) [129] unifies the advantages of automatic differentiation and numerical differentiation, providing more robust and efficient training than automatic differentiation-based PINNs.
Fifth, in order to improve the generalization of PINNs, a transfer PINN [131] is used to solve forward and inverse problems in nonlinear partial differential equations by parameter sharing. Sixth, PINNs struggle in cases where the target functions to be approximated exhibit high-frequency or multi-scale features. A sea of variants of PINNs are proposed to address this problem. For example, finite basis PINNs [132] use ideas from domain decomposition for complex, multi-scale solutions. Krishnapriyan et al. [133] use curriculum PINN regularization and sequence-to-sequence learning to address the problems of PINNs that fail to learn relevant physical phenomena for even slightly more complex problems. In addition, spatio-temporal and multiscale random Fourier features in multiscale PINNs [134] are used to address the high-frequency or multi-scale problems in PINNs. Seventh, adaptive algorithms for weighing components of the PINN loss function (hyperparameter selection) have also been proposed, such as self-adaptation PINNs [135], a heuristic method [136].
Finally, PINNs are used to solve the inverse problem of parameter identification (data-physics-driven parameter discovery). Inverse problems to parameters’ estimation are in general ill-posed because the problem is non-unique due to the high number of unknowns. Furthermore, the stability of the solution to noise in the data and modeling errors is generally not guaranteed. The physics-informed neural network theory of functional connections (PINN-TFC) [137], for example, the extreme theory of functional connections, mitigates these issues. A PINN with sparse regression [152] is proposed to discover governing partial differential equations from scarce and noisy data for nonlinear spatiotemporal systems. This discovery approach seamlessly integrates the strengths of NNs for rich representation learning, physics embedding, automatic differentiation, and sparse regression. Saqlain et al. [138] select a sequence of progressively more complex yet physically relevant examples and seek to leverage the PINN in an inherently discrete, high-dimensional setting. Addressing such inverse problems through the use of PINNs is still at an early stage, and further studies are certainly warranted.
2.2.2 Physics-informed Neural Operators
To overcome the challenges of vanilla PINNs, physics-informed neural operators (PINOs) [140] are proposed. These utilize both the data and equation constraints (whichever are available) for deep operator networks [153]. One particular PINO, physics-informed DeepONet [139], trains the DeepONet by directly incorporating known differential equations into the loss function, along with some labeled datasets. The outputs of the DeepONet are differentiable with respect to their input coordinates, thereby allowing the use of automatic differentiation to develop an appropriate regularization mechanism for biasing the target output functions to satisfy the underlying PDE constraints. However, the PDE loss of physics-informed DeepONet is computed at any query point. The input sensors limit this PINO to a fixed grid or basis, and therefore it is not discretization invariant. In addition, its architecture comes with the limitations of linear approximation. Furthermore, in order to enhance the computational efficiency and generalization of physics-informed DeepONet, a subset of problem-specific surrogate models based on the physics-informed DeepONet is designed, such as physics-informed DeepONet for chemical kinetics [154], and physics-informed variational formulation of DeepONet for brittle fracture analysis [18]. Furthermore, in order to achieve discretization invariance, the physics-informed FNO (PINO) is proposed [140] as an integration of operator learning and physics-informed settings. PINO reduces the labeled dataset requirement for training the neural operator and helps reach convergence of the solution faster. In this setting, it is important to note that, unlike DeepONet, FNO outputs the solution on a grid that employs spectral methods, finite difference, or other numerical gradient solution methods. However, PINO has limited accuracy in solving non-periodic problems due to the numerical gradient method and relies on certain training samples to achieve better accuracy. In addition, the adaptability to different PDE solution problems is not strong due to the selection of model parameters. It cannot be applied to large-scale and long-term sequence PDE solutions. In order to address the problem that PINO cannot represent nonperiodic functions, an architecture that leverages Fourier continuation (FC-PINO) [141] is proposed to apply the exact gradient method to PINO. In addition, a geometry-adaptive physics-informed neural solver [142] is proposed based on the advantages of no training data in physics-informed neural networks; this solver also possesses a fast, accurate, geometry-adaptive and resolution-invariant architecture. Most recently, a physics-informed WNO [143] is proposed for learning the solution operators of families of parametric PDEs without labeled training data, based on the advantage of time-frequency localization of wavelets.
2.3 Physics-embedded Machine Learning
PeML is achieved by embedding physics in the model frameworks or modules. It can be divided into physical equation-embedded ML (PEeML), physical property-embedded ML (PPeML), and physical condition-embedded ML (PCeML).
2.3.1 Physical Equation-embedded Machine Learning
PEeML uses the knowledge of specific equations (such as differential equations) to design the frameworks or modules of machine learning. The objectives and limitations of different PEeML are briefly discussed in Table 3.
Specifically, a simple, yet powerful and general framework (AutoIP) [155] for automatically incorporating physics that can integrate all kinds of differential equations into Gaussian processes to enhance prediction accuracy and uncertainty quantification is proposed. Physics-aware finite volume neural network (FINN) [156] is introduced for learning spatio-temporal advection-diffusion processes by modeling the constituents of PDEs in a compositional manner. Deep Lagrangian networks (DeLaN) [157] encode the Euler-Lagrange equation originating from Lagrangian mechanics. DeLaN can be trained using standard end-to-end optimization techniques while maintaining physical plausibility. A framework called MeshGraphNets [158] is proposed to solve the underlying PDEs using graph neural networks. MeshGraphNets works by encoding the simulation state into a graph and performing computations in two separate spaces: the mesh space, spanned by the simulation mesh, and the Euclidean world space in which the simulation manifold is embedded. By passing messages in mesh space, we can approximate differential operators that underpin the internal dynamics of most physical systems. Furthermore, fluid graph networks (FGN) [159] using graphs to represent the fluid field, is proposed for solving a Lagrangian representation of the Navier-Stokes equation. FGN is able to retain many important physical properties of incompressible fluids, such as low-velocity divergence. Several follow-up works are introduced using graph neural networks [160, 161]. Apart from building an explicit graph structure, an efficient ConvNet architecture based on a continuous convolution layer [71] is developed. The network processes sets of particles in which dynamic particles and static particles are used to represent the fluid and describe the boundary of the scene, respectively.
Furthermore, the use of differentiable modules in NNs is a promising direction for physical equation-embedded ML. It implements physical equations as differentiable feature space or computational graphs [162], enabling the optimization of dynamic processes with analytical gradients and therefore improving sample efficiency. For example, a differentiable physics-informed graph network (DPGN) [163] is proposed to incorporate implicit physics equations in latent space. A differentiable layer [164], called PDE-Constrained-Layer, is developed by using implicit differentiation, thereby allowing us to train NNs with gradient-based optimization methods. A spatial difference layer [165] is utilized in physics-aware difference graph networks (PA-DGN) to efficiently exploit neighboring information under the limitation of sparsely observable points. A physics-based finite difference convolution connection [166] in a physics-encoded recurrent CNN is introduced to facilitate the learning of the spatiotemporal dynamics in sparse data regimes. Furthermore, the term differentiable modeling is proposed to include any method that can produce gradients rapidly and accurately at scale, potentially serving as the basis for unifying NN and process-based geoscientific modeling [13].
Table 3: Proposed Physical Equation-embedded Machine Learning
Methods
Physical equations
Embedding methods and objectives
Limitation(s)
AutoIP [155]
Differential equations
Automatically incorporating physics that can integrate all kinds of differential equations into Gaussian processes
Limitations on large-scale applications
FINN [156]
Advection-diffusion type equations
Learning spatio-temporal advection-diffusion processes by modeling the constituents of PDEs in a compositional manner (the incorporation of explicitly defined physical equations)
Limited scalability (beyond second-order spatial derivatives); limitation on the real-world applications
DeLaN [157]
Euler-Lagrange equation from Lagrangian mechanics
Deep Lagrangian networks encoding the Euler-Lagrange equation originating, which can be trained using standard end-to-end optimization techniques while maintaining physical plausibility.
Limited representational power of the physics prior when applied to a physical system with complex dynamics
MeshGraphNets [158]
Mesh-based partial differential equations
A framework for learning mesh-based simulations using graph neural networks: the mesh-space, spanned by the simulation mesh, and the Euclidean world-space in which the simulation manifold is embedded.
High computational complexity; hyperparameters
FGN [159]
A Lagrangian viewpoint of the Navier-Stokes equation
Fluid graph networks (FGN) uses graphs to represent the fluid field. In FGN, fluid particles are represented as nodes and their interactions are represented as edges. FGN decomposes the simulation scheme into separate parts, including advection, collision, and pressure projection
Limited generalization and computational efficiency
Continuous convolutions [71]
Lagrangian fluid simulation of the Navier-Stokes equation
An efficient ConvNet architecture, based on continuous convolution layer, processes sets of particles. Dynamic particles and static particles are used to represent the fluid and describe the boundary of the scene, respectively
Limited computational efficiency; cannot deal with deformable solids
Differentiable modules
Differential equations
These modules implement physical equations as differentiable feature space or computational graphs, enabling the optimization of dynamic processes with analytical gradients and therefore improving sample efficiency. Examples include differentiable physics-informed graph networks [163], a differentiable layer (PDE-Constrained-Layer) [164], physics-aware difference graph networks [165], physics-based finite difference convolution connection [166], and differentiable modeling [13]
Limitations on large-scale applications; require an in-depth understanding of differential equations.
2.3.2 Physical Property-embedded Machine Learning
Physical properties usually include symmetry and conservation of physical systems. By building a neural network that inherently respects a given physical property, we thus make conservation of the associated quantity more likely and consequently the model’s prediction more physically accurate.
Conservation. Many real-world systems are governed by conservation laws related to mass, energy, momentum, charge, or particle counts, which are often expressed through continuity equations. For example, in hydrology, it is the amount of water [167]. However, standard deep learning approaches (such as CNNs and LSTM) struggle at conserving quantities across layers or timesteps. Thus, mass-conserving LSTM (MC-LSTM) [168] adheres to these conservation laws by extending the inductive bias of LSTM to model the redistribution of those stored quantities. MC-LSTMs [168] modify the recurrence in traditional LSTM to guarantee the conservation of the mass input. The key idea is to use the memory cells from LSTMs as mass accumulators, or mass storage. In addition, an antisymmetrical continuous convolutional layer [169] in a hierarchical network is presented by enforcing the conservation of momentum with a hard constraint. The parameterization of deep neural networks [170] is designed to satisfy the continuity equation, a fundamental conservation law, by building divergence-free neural networks. An implicit neural network layer [171] that incorporates the conservation of mass is proposed by using implicit differentiation. Furthermore, a systematic way of enforcing nonlinear analytic constraints (energy conservation and mass conservation) in neural networks via constraints in the architecture [172] is also proposed. The concept of “conversion layers” is introduced in the architecture that transforms nonlinearly constrained mappings into linearly-constrained mappings within NNs without overly degrading performance.
Symmetry has long been implicitly used in DL to design networks with known invariances and equivariances. Let $f:X\rightarrow Y$ be a function and $G$ be a group. Assume $G$ acts on $X$ and $Y$. If $f(gx)=gf(x)$ for all $x\in X$ and $g\in G$, the function $f$ is $G$-equivariant. If $f(gx)=f(x)$ for all $x\in X$ and $g\in G$, the function $f$ is G-invariant.
Convolutional neural networks enabled breakthroughs in computer vision by leveraging translation equivariance. Similarly, recurrent neural networks, graph neural networks, and capsule networks all impose symmetries [12]. There is a deep connection between symmetries and physics. For example, Noether’s law gives a correspondence between conserved quantities and groups of symmetries [173]. In addition, the symmetries of fluid dynamics [174] are exploited to design equivariant networks. The Navier-Stokes equations and heat equations are invariant under the following transformations:
- Space translation: $T_{c}^{\mathrm{sp}}w(x,t)=w(x-c,t),\quad c\in\mathbb{R}^{2}$,
- Time translation: $T_{\tau}^{\text{time}}w(x,t)=w(x,t-\tau),\quad\tau\in\mathbb{R}$,
- Uniform motion: $T_{c}^{\mathrm{um}}w(x,t)=w(x,t)+c,\quad c\in\mathbb{R}^{2}$,
- Rotation/Reflection: $T_{R}^{\text{rot }}w(x,t)=Rw\left(R^{-1}x,t\right),R\in O(2)$,
- Scaling: $T_{\lambda}^{sc}w(x,t)=\lambda w\left(\lambda x,\lambda^{2}t\right),\quad\lambda\in\mathbb{R}_{>0}$.
Recently, incorporating symmetries of physics into NNs has been developed in the context of CNNs [175], GNNs, and neural operators [176].
First, an end-to-end ML-based deep potential–smooth edition (DeepPot-SE) [177] is constructed. DeepPot-SE is extensive and continuously differentiable, scales linearly with system size, and preserves all the natural symmetries of the system. Translational, rotational, and permutational symmetries are preserved by symmetry preserving functions.
Second, there have been a number of GNN-based works, called geometrically equivariant graph neural networks [178], that leverage symmetry as an inductive bias in learning to simulate. These models are designed such that their outputs will rotate/translate/reflect in the same way as the inputs, hence retaining the symmetry. For example, EGNN [179] models the interaction with the invariant distance as input, computed as an inner product of the relative positions, and GMN [180] generalizes to a multi-channel version by taking into consideration a stack of multiple vectors. Subequivariant GNN [181] relaxes equivariance to subequivariance by considering external fields like gravity, where the universal approximation ability holds theoretically.
Third, an integral neural operator architecture [176] is designed to learn physical models with fundamental conservation laws automatically guaranteed. The translation- and rotation-invariant neural operator is developed by replacing the frame-dependent position information with its invariant counterpart in the kernel space; consequently it abides by the conservation laws of linear and angular momentums.
Causality. A fundamental pursuit in physics is to identify causal relationships. Integrating causality into ML promises to advance the comprehension of physical processes and bolster the robustness of ML models [182]. For example, causality-DeepONet [183] implements the physical causality in the DeepONet structure. Specifically, the causality-DeepONet is applied to learn the operator representing the response of a building due to earthquake ground accelerations. However, many questions of causality-embedded ML remain open, such as using causality to improve deep learning models and understand the system’s response under interventions, disentangling complex and multiple treatments, and designing the environment to control the given physics. In addition, a causal training algorithm  [184] is introduced to restore physical causality during the training of a PINNs model by appropriately re-weighting the PDE residual loss at each iteration of gradient descent. This simple modification provides PINNs with the capability to address complex problems, such as the incompressible Navier-Stokes equations in the turbulent regime.
2.3.3 Physical Condition-embedded Machine Learning
There are many physical conditions in the physical system, such as boundary conditions and initial conditions in PDEs, boundaries of land use and land cover in the physical environment, and so on. Violating these conditions can lead to unstable models and non-physical solutions. Thus, it is critical for ML to satisfy these constraints in order to capture the underlying physics accurately and provide reliable models for rigorous research and engineering design. Specifically, for PDEs, three types of boundary conditions (BCs) that are widely used to model physical phenomena are (1) Dirichlet: fixed-value at the boundary; (2) Neumann: fixed-derivative at the boundary; and (3) periodic: equal values at the boundary. In order to embed these BCs into NNs, a boundary enforcing operator network [185] is proposed that enables the BC satisfaction of neural operators by making structural changes to the operator kernel. The physics-embedded neural network [186] is developed by considering BCs in GNN.
2.4 Physics-aware Hybrid learning
PaHL directly combines pure physics-based models, such as numerical methods, climate, land, hydrology, and Earth system models, with ML models. As shown in Fig. 4, according to the hybrid way, hybrid learning can be divided into serial, parallel, and complex ways of approaching learning.
Figure 4: Different hybrid approaches in physics-aware hybrid learning.
2.4.1 Neural–numerical Hybrid Learning
PDEs are notoriously difficult to solve, and traditional numerical approximation schemes have high computational costs. Recently, hybrid neural-numerical solvers have been developed to complement the modern trend of fully end-to-end learning systems.
Serial Hybrid has broad potential advantages for physical models and data-driven models. For example, NNs are used to handle data coming from irregular domains. A family of neural ODEs [187, 188] changes the traditionally discretized neuron layer depths into continuous equivalents such that the derivative of the hidden state can be parameterized using a neural network. The output of the network is then computed using a black box differential equation solver, making Neural ODEs an efficient combination of neural nets and numerical solvers. In addition, the advantages of physics solvers (such as their scale-invariant properties) can be imparted into NNs. A novel hybrid training approach [189] that combines a scale-invariant physics solver with ML techniques is developed for solving inverse problems, such as parameter estimation and optimal control. Gradients of NNs are replaced by the update computed from a higher-order solver that can encode the scale-invariant property of the physics.
Parallel Hybrid methods are designed to develop fast, robust, and reliable solvers for solving PDEs. Physics-driven models are usually used as a supervised signal to train data-driven NNs. For example, a hybrid, iterative, numerical, and transferable solver [190] called HINTS is proposed for differential equations. HINTS combines standard relaxation methods (such as Jacobi, Gauss-Seidel, conjugate gradient, multigrid methods, and their more advanced variants) and DeepONet. Compared to standard numerical solvers, HINTS is capable of providing faster solutions for a wide class of differential equations, while preserving accuracy close to machine zero. In addition, a hybrid graph neural network [191] that combines a traditional graph convolutional network with an embedded differentiable fluid dynamics simulator is developed. The hybrid graph neural network both generalizes well to new situations and benefits from the substantial speedup of neural network CFD predictions, while also substantially outperforming the coarse CFD simulation alone. A fast, differentiable hybrid approach [192] combines a 2D direct numerical simulation for the deformable solid structure and a physics-constrained neural network surrogate to capture the hydrodynamic effects of the fluid.
Complex Hybrid methods are flexible frameworks that can be used to address many problems in solving PDEs. For example, in order to address the scalability issue of solving PDEs on a large scale, a physics-aware downsampling method [193] is proposed by minimizing the distance between the solutions on the fine and coarse grids. Specifically, a serial hybrid method is designed to downsample fine grid terrain, by combining a numerical solver (such as finite difference) on coarse grid terrain and a downsampling neural network. Then, a parallel numerical solver on fine grid terrain is used to provide a supervised signal for the predictions on the coarse grids.
2.4.2 Neural–hydrology Hybrid Learning
The hydrological model, a standard representation of physical processes, serves as an input-output model simulating the evolution and dynamics of water storage, fluxes, and physical properties in the Earth’s surface and subsurface [2]. Neural-hydrology hybrid learning utilizes ML methods to harness and integrate a broad variety of hydrological models into a final process-based prediction product. Recently, neural–hydrology hybrid learning is receiving growing attention due to interpretability in hydrological models, a better appreciation of ML, and the availability of computing resources and methods [30]. Here we list some recent neural–hydrology hybrid methods for hydrological variable prediction, such as streamflow and runoff.
Serial Hybrid combines the hydrological models and data-driven models sequentially. Serial hybrid includes a conceptual hydrological model that seeks to combine the strengths of data-driven and conceptual models, driven by dynamic predictions. For example, a hybrid monthly streamflow forecasting approach [194] is explored, where simulated soil moisture from the GR4J conceptual rainfall-runoff model is used to represent initial catchment conditions in a Bayesian NN model. In this work, the hybrid model performed better than either the GR4J model or the Bayesian neural network alone. In addition, to reduce the computational complexity of parallel hybridization of the conceptual rainfall-runoff models and ML techniques, NNs and support vector regression are embedded into a monthly lumped conceptual rainfall-runoff model [195]. Mohammadi et al. [196] use two process-driven conceptual rainfall-runoff models (HBV: Hydrologiska Byrans Vattenbalansavdelning, and NRECA: Non Recorded Catchment Areas) to provide inputs to support vector machines (SVMs) and an adaptive neuro-fuzzy inference system (ANFIS), in order to build seven variants of hybrid models. Results showed that AI-based hybrid models generally led to more accurate streamflow estimates compared with HBV and NRECA. In addition, three hybridization approaches that combine the simulations of the HBV model with LSTM [197] are developed for semi-arid regions. The results show distinct improvements in the three types of hybrid models when compared with the HBV model and standalone LSTM model. A number of studies have coupled conceptual models and data-driven models with serial hybrid methods.
Parallel Hybrid. In a parallel hybrid structure, the data-driven and dynamic models are combined in parallel. Parallel hybrid models can employ a data-driven model to combine hydrological predictions in parallel. For example, a combination of conceptual and ML techniques [198] is proposed for enhancing the accuracy of streamflow simulation in a snow-covered basin. An end-to-end hybrid modeling approach [199] is presented to learn and predict spatial-temporal variations of observed and unobserved hydrological variables globally. The model combines a dynamic neural network and a conceptual water balance model, constrained by the water cycle observational products of evapotranspiration, runoff, snow-water equivalent, and terrestrial water storage variations. The model reproduces observed water cycle variations very well and the emergent relations of runoff-generating processes are qualitatively consistent with our understanding. Furthermore, the hybrid hydrological model [200] is used for global hydrological modeling that exploits the data adaptivity of neural networks for representing uncertain processes within a model structure based on physical principles (e.g., mass conservation). The hybrid model is trained simultaneously with a multi-task learning approach. It can conclude that the hybrid model provides a new data-driven perspective on modeling the global hydrological cycle and physical responses with machine-learned parameters that is consistent with and complementary to existing global modeling frameworks. In addition, a parallel hybrid approach can be employed to correct model biases [201].
Complex Hybrid. In a complex structure, data-driven models can be integrated into the input space, output space, and interior of hydrological models in a complex hybrid method, according to different application requirements. Specifically, a general physics-AI approach [9] is proposed to improve AI geoscientific awareness, wherein temporal dynamic geoscientific models are included as a special process-wrapped recurrent neural network (denoted as P-RNN). In addition, the common NN layers without physics address processes unrepresented by the P-RNN layer. The experiments of the physics-AI approach demonstrate that the complex hybrid model has enhanced prediction accuracy, robust transferability, and good intelligence for inferring unobserved processes. Furthermore, a differentiable programming framework [202] is designed to parameterize, enhance, or replace the process-based model’s modules in a simple hydrologic model, HBV, for predictions of hydrologic variables (streamflow). Specifically, an LSTM outputs the physical parameters for a process-based model. In process-based models, certain parts of the model can be replaced by neural networks and the structure can be updated. The framework is trained in an end-to-end fashion and there are no intermediate supervising data (or labels) for the outputs of the neural networks. The main loss function is calculated between the process-based model output and observations.
2.5 Summary of Physics-aware Machine Learning
These four methods in the PaML community, including their corresponding benefits and drawbacks for scientific problems, are summarized in Table 4.
Table 4: Four Approaches to Physics-aware Machine Learning
Term Description Methods Benefits and drawbacks
Physical Data-guided Machine Learning (PDgML)
PDgML is a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data.
Classical physical data-guided neural networks (spatio-temporal modeling, dynamic system modeling), deep operator networks, physics-discovery neural networks
[Benefits] Only need simulations from physics model, fast inference, the discovery of the dynamic behavior;
[Drawbacks] The architecture and learning process are black box, limited generalization and transferability.
Physics-informed Machine Learning (PiML)
PiML is a widely used approach to incorporating physical constraints that can be trained from additional information obtained by enforcing the physical laws (for example, designing loss functions or regularization).
Physics-informed neural networks (PINNs), physics-informed deep neural operators (PINOs)
[Benefits] Incomplete models and imperfect data, strong generalization in small data regime, understanding deep learning by the optimization process, tackling high dimensionality; uncertainty quantification, data-physics-driven parameters discovery;
[Drawbacks] Effectiveness and adaptability, data generation and benchmarks, large-scale applications, the architecture and the learning process are still black boxes.
Physics-embedded Machine Learning (PeML)
PeML is achieved by embedding physics in the model frameworks or modules.
Physical equation-embedded machine learning (PEeML), physical property-embedded machine learning (PPeML), physical condition-embedded machine learning (PCeML)
[Benefits] Interpretability of architecture and learning process, generalization;
[Drawbacks] Computational cost is relatively high, hard inductive biases from physics may hurt the representation power of neural nets, and require an in-depth understanding of the physical principles to design.
Physics-aware Hybrid Learning (PaHL)
PaHL directly combines pure physics-based models, such as numerical methods, climate, land, hydrology, and Earth system models with ML models. Depending on the hybrid approach, hybrid learning can be categorized as serial, parallel, or complex.
Neural–numerical hybrid learning, neural–hydrology hybrid learning
[Benefits] Can leverage both the flexibility of data-driven models and the interpretability and generalizability of physics-driven models, speed, and operational convenience;
[Drawbacks] Need an optimal balance between the physics-driven and data-driven models.
3 HydroPML: Process-based Hydrology in Physics-aware Machine Learning
As shown in Fig. 5, the proposed HydroPML includes hydrodynamic process understanding and rainfall-runoff hydrological process understanding.
Figure 5: The proposed process-based hydrology in physics-aware machine learning (HydroPML). Red labels denote different objectives of HydroPML. For example, the oval labeled 1.1 represents short-term forecasts. See the text for more detail.
Figure 6: The operational flood forecasting system based on HydroPML
3.1 Rainfall-runoff Process Understanding
Rainfall-runoff models are extensively utilized in hydrology to investigate hydrological processes and play a crucial role in water resources management, encompassing flood prediction [203] and drought analysis [204]. These models have been developed for diverse applications, ranging from small catchments to global scales.
3.1.1 Physical Methods in Hydrological Process
The hydrological model can be classified into three broad groups: the empirical model, conceptual models, and physical-based models [2]. Empirical models only take information from the existing data without considering the features and processes of hydrological systems and hence these models are also called data-driven models. Physical and conceptual models need a thorough understanding of the physics involved in the movement of surface water in the hydrological cycle [205].
Conceptual models interpret runoff processes by connecting simplified components in the overall hydrological process with the help of a number of interconnected storages used to represent different components. The conceptual models are usually lumped in nature and use the same value of parameters for the whole watershed. These models need a range of parameters and meteorological input data. A large number of meteorological and hydrological records is required for parameter calibration. A host of conceptual models have been proposed in the past including HBV model [206], non-recorded catchment areas (NRECA) [207], Boughton model [208], ARNO model [209], Xinanjiang model [210], etc.
Physical-based models are based on an understanding of the physics related to hydrological processes. The general physics laws and principles used include water balance equations, conservation of mass and energy, momentum, and kinematics. Finite difference or finite computation schemes are generally used to solve these equations. For being built on physical laws and equations, spatially-explicit, process-based models are, in theory, capable to respond correctly to dynamic changes in catchment conditions and climate forcing; they can thus be used for the assessment of non-linearities between states, drivers, and catchment responses [211, 212]. Therefore, they are practical tools to predict, project, and attribute impacts of global change (e.g. climate change, land use change, water management) on the water cycle and even hydrological extremes. The greatest strength of a physical model is the connection between model parameters and physical catchment characteristics, which makes it more realistic. As physical models are watershed semi- or fully distributed in nature, they require a large amount of data such as soil-hydraulic properties, vegetation, land cover characteristics, initial water depth, topography, or dimensions of the river network to explore the changes in the hydrological cycle due to human interference and climate change for water resources management [213]. Some examples of physical models include soil and water assessment tool (SWAT) [214], system hydrologique european (SHE) [215] or MIKE SHE [216], visualizing ecosystem land management assessments (VELMA) [217], kinematic runoff and erosion model (KINEROS) [218], Penn State integrated hydrologic modeling system (PIHM) [219], variable infiltration capacity (VIC) [220], and others.
3.1.2 Rainfall-runoff Forecast Meets Physics-aware Machine Learning
The primary input to the rainfall-runoff models is past forcing data (such as precipitation, temperature), constant regional attributions (such as land use land cover (LULC), topography, and others) and the output is the future runoff or streamflow. As shown in Table 5, improving process-based rainfall-runoff models requires progress on several fundamental research challenges: (1) building appropriate methods to forecast the runoff for the different time scales within the model domain (Short-term Forecasts and Long-term Forecasts); (2) representing the variability of hydrologic processes across a hierarchy of spatial scales (Spatial Variability); (3) verifying model reliability (Bias Reduction and Model Reliability); (4) testing a model across the different model subdomains or ungauged basins (Missing Data and Ungauged Basins); (5) estimating input data and model parameters (Parameterization); and (6) characterizing model uncertainty (Uncertainty Estimation in Rainfall-runoff Forecast).
Table 5: PaML-based Rainfall-runoff Hydrological Forecasts Classified by Objectives
Objectives PDgML PiML PeML PaHL
(1.1) Short-term Forecasts
LSTM [221, 222], physics-guided LSTM [43], GRUs [223], graph convolutional GRUs [224], graph convolutional RNN [225], stacked LSTM and bi-directional LSTM [226], ensemble model [227, 228], ConvLSTM [229], data-driven forecasting using only in-stream measurements [230]
GW-PINN [231]
Hybrid method combining hydrologic processes with LSTM [232], SWAT-LSTM [233], conceptual hydrological model by integrating NNs [234], urban rainfall-runoff modeling [235], Google’s operational flood forecasting system [236]
(1.2) Long-term Forecasts
LSTM [10], Gaussian process regression [237], deep belief network [238], extreme learning machine [239], an encoder-decoder CNN algorithm [240], Bayesian ensemble learning by combining different ML methods [241]
LSTM by considering conservation of mass into loss function [15]
Entity-aware-LSTM [242], MC-LSTM [168], differentiable modeling [13], P-RNN layer [9]
A hybrid monthly streamflow forecasting approach [194], ML is embedded into a lumped conceptual rainfall-runoff model [195], three hybridization approaches [197], differentiable parameter learning [243]
(1.3) Spatial Variability
Fully-distributed GNN [244], graph convolutional GRUs based model [224], graph convolutional RNN [225], ConvLSTM [229], graph-based reinforcement learning [245], data synergy [246], initializing a recurrent GNN by physical model [247]
HydroNets [248], a differentiable, learnable physics-based routing model [249]
Hybrid approach combining a spatially-distributed snow model with convLSTM [250]
(1.4) Bias Reduction and Reliability
Bias-corrected remote sensing products [251], directly measured datasets [230], ensemble methods [252, 253]
MC-LSTM [168, 254], differentiable modeling [13]
LSTM using an additional input feature from the process-based model [255], using physical models coupled with LSTM [256], WRF-Hydro-LSTM [257], LSTM daily streamflow prediction models [258]
(1.5) Missing Data and Ungauged Basins
Attention-based deep learning (such as 3D-CNN-Transformer) [259], input-selection ensemble method [260]
AI-based hybrid models [196], serial hybrid LSTMs [261, 262], transfer learning in SWAT-LSTM [233].
(1.6) Parameterization
A machine learning approach [263], deep neural network [264], machine learning-based maps [265]
A differentiable neural network [202], differentiable parameter learning [243]
(1.7) Uncertainty Estimation in Rainfall-runoff Forecast
Data-driven approach [266], Bayesian processor [267], hybrid ensemble and variational data assimilation framework [268], hierarchical model tree [269], probabilistic deep ensembles [270], LSTM-based approaches [271, 272], uncertainty estimation benchmarking procedure and baselines [273], analysis framework for sources of uncertainties [274]
Using physical models coupled with LSTM [256], XAJ-MCQRNN [275], hybrid error correction model [276]
Short-term Forecasts. See Fig. 5 (label 1.1). Short-term forecasts focus on outlook horizons of hours to weeks. Innovative forecasting tools will reform river discharge forecasting, flood early warning mechanisms, and agricultural and hydropower management schemes. PaML offers robust tools to address the formidable challenges associated with the uncertainties, data dependencies, and intricacies inherent in short-term hydrological forecasting. First, the high uncertainty and complexity associated with hydrological processes and weather-climate factors affecting river basins have increasingly led researchers to use PDgML approaches. Specifically, the PDgML models based on LSTM [221, 222] are proposed to predict streamflow. They have been found to have greater accuracy in predicting hourly and daily streamflow than a classical neural network with back-propagation. A model [223] that uses multiple GRUs is proposed to predict up to 120 hours of streamflow. Granata et al. [227] show how the ensemble model based on random forest and multilayer perceptron outperforms the bi-directional LSTM network in predicting peaks of flow rates, as well as having significantly shorter computation times. Wegayehu and Muluneh [226] also compare different ML algorithms for one-step daily streamflow forecasting and show how both MLP and GRU algorithms outperform stacked LSTM and bi-directional LSTM on a nearly equal basis. An ensemble model [228] based on the combination of the nonlinear autoregressive network, multilayer perceptron, and random forest is proposed for short-term streamflow forecasting, considering precipitation as the only exogenous input and a forecast horizon of up to 7 days. In addition, to model the spatial diversity of rivers, the graph convolutional GRUs-based model [224] is developed to predict the next 36 hours of streamflow for a sensor location using the upstream river network. Graph convolutional RNN [225] is developed for robust water demand forecasting. More detailed information about PDgML studies on short-term streamflow prediction can be found in [277]. Furthermore, to overcome data dependence, attempts have been made based on the PiML community, such as PINNs, for solving groundwater flow equations [231]. Because it is difficult to express the whole process of rainfall-runoff with specific equations, the current progress based on PiML technology is very slow. In addition, the hybrid method combining hydrologic processes with LSTM [232] is successful in assimilating recent observations of streamflow to improve near-term daily forecasts of streamflow. A coupled soil and water assessment tool (SWAT)-LSTM approach [233] is developed to provide a potential shortcut to achieving daily streamflow simulations in both ungauged and poorly gauged watersheds. The conceptual hydrological model by integrating NNs [234] reveals that the hybrid model has better prediction performance than either the original GR4J or the single NN–based runoff prediction model.
In general, existing physical and conceptual models are well-established in operational forecasting within the time scale of hours to weeks. The development of PaML-based short-term forecast methods is not very mature [30]. However, PaML offers promising alternatives to physics-based hydrological and inundation models for flood forecasting due to its automation, efficiency, and flexibility. For example, the PDgML (ConvLSTM [229], data-driven forecasting using only in-stream measurements [230]) for extracting spatio-temporal features of hydrological information outperforms the recent models in terms of flood arrival time and peak discharge. PaHL-based methods (such as Google’s operational flood forecasting system [236] and urban rainfall-runoff modeling [235]) show potential for flood nowcasting.
Long-term Forecasts. See Fig. 5 (label 1.2). Long-term forecasts focus on outlook horizons of sub-seasonal, year, to decade. The vast majority of ML predictions in the literature are PDgML methods, which are driven by observational or historical data. PDgML allows us to capture complex non-linear patterns from data that would otherwise require extrapolation, thus providing an alternative approach to long-term rainfall-runoff forecasts, for example, LSTM [10], Gaussian process regression [237], deep belief network [238], extreme learning machine [239], an encoder-decoder CNN algorithm [240], and Bayesian ensemble learning by combining different ML methods [241]. These methods have good accuracy in the study of the rainfall-runoff relationship and climate factor-rainfall relationship. In order to alleviate the dependence on data, a physics-guided LSTM [15] that considers the conservation of mass into loss function is proposed. Digging deeper into physics (such as mass conservation and similarities) in the hydrological process, MC-LSTM [168] is proposed by modifying the LSTM to guarantee the conservation of the mass input. Entity-Aware-LSTM [242] is presented for learning catchment similarities as a feature layer in a deep learning model. In addition, based on the PeML technology, some NN modules with physical properties (such as differentiable modeling [13] and the P-RNN layer in the general physics-AI approach [9]) are proposed for rainfall-runoff forecasts. Furthermore, hybrid learning models (PaHL) can be suitable alternatives to hydrological models. Hybrid learning based on conceptual hydrological models tries to combine the strengths of hydrological models and data-driven methods. For example, a hybrid monthly streamflow forecasting approach [194] and three hybridization approaches [197] show that the hybrid model is not only significantly more precise than conceptual hydrological models but also outperforms a purely ML model.
However, in most hybrid modeling frameworks, variables computed by the conceptual model are used as inputs to a data-driven model, which necessarily increases computation time. In addition, those interactions often go unaccounted for because the two models are calibrated separately. Thus, embedded ML in a lumped conceptual rainfall-runoff model [195] is proposed by coupling the two models and performing their calibration jointly, and differentiable parameter learning [243] is presented to efficiently learn a hybrid model in an end-to-end manner.
Spatial Variability. See Fig. 5 (label 1.3). Rainfall-runoff modeling is difficult due to the non-linear structure of the process, which involves much spatio-temporal variability. Spatial variability is critical for predictions in ungauged basins and for developing universal regional models. PaML (especially GNNs) provides the possibility of distributed modeling. Specifically, for PDgML, GNN [244] is used for fully-distributed rainfall-runoff modeling. The result shows that the GNN-based fully-distributed model [244] successfully represents the spatial information in predictions. A graph convolutional GRUs based model [224], ConvLSTM [229], and graph convolutional RNN [225] can also be developed to model spatial diversity of rivers. A recurrent GNN [247] is proposed to capture the interactions among multiple segments in the river network. In this work, a physics-based model is used to initialize the machine learning model and learn the physics of streamflow. To reduce the need for large training datasets, a real-time active learning method that uses spatial and temporal contextual information to select representative query samples is employed in a graph-based reinforcement learning framework [245]. The effectiveness of graph-based reinforcement learning is demonstrated by predicting streamflow and water temperature in the Delaware River basin. In addition, region-specific differences can be learned for the LSTM model by inputting sufficient heterogeneous data (data synergy) [246].
Furthermore, spatial physical characteristics of rivers (such as river networks and hydrologic routing modules) can be encoded into NNs based on the PeML technology. For example, a hydrologic model called HydroNets [248] is built by leveraging river network structure. Specifically, each node in the graph represents a basin, and an edge direction corresponds to water flow from a sub-basin to its containing basin. A differentiable, learnable physics-based routing model [249] is proposed to constrain spatially-distributed parameterization. This is the first time an interpretable, physics-based model is learned on the river network to infer spatially-distributed parameters. Finally, for the PaHL, the hybrid models combining conceptual models and data-driven models are often lumped, i.e., spatially averaged, at the catchment scale. To address this problem, a hybrid approach [250] is developed to predict streamflow. The hybrid approach combines a spatially-distributed snow model with convLSTM, which simulates the effect of surface and subsurface properties on the streamflow. In the future, building a robust spatio-temporal representation for different watersheds based on the PaML community will be an ongoing direction for research.
Bias Reduction and Reliability. See Fig. 5 (label 1.4). One important challenge of PaML-based rainfall-runoff forecasts is the ability of such models to produce reliable or physically-plausible results in different domains, especially extreme conditions such as floods and droughts. Specifically, reliable spatio-temporal remote sensing data and directly measured datasets can be used to reduce the bias of PDgML-based rainfall-runoff forecasting. For example, bias-corrected remote-sensing precipitation products are adopted as precipitation input of a spatio-temporal DL rainfall-runoff forecasting model [251], which can improve the model’s reliability for extreme flood forecasts. Directly measured datasets are utilized for streamflow estimation to develop a data-driven forecasting algorithm [230]. However, these bias-corrected data-based approaches suffer from deficiencies over small catchments due to uncertainty in remote sensing data and errors in hydrological models. Thus, some ensemble methods (such as resampling ensemble methods [252], which combine hydrometeorological modeling and LSTM [253]) are explored to improve the reliability and robustness of PDgML models. Furthermore, one emerging route for bias reduction and reliability is to employ PeML designs that explicitly observe the physical laws in the rainfall-runoff process, such as MC-LSTM [168] and differentiable modeling [13]. Another new development is hybrid models to produce physically-plausible or explainable results. For instance, Wi et al. [255] examine whether deep learning models can produce reliable future projections of streamflow under warming. It concludes that the hybrid LSTM using estimates of evapotranspiration from the process-based model as an additional input feature may support the use of LSTMs for hydrologic projections under climate change, but so may training LSTMs to a large, diverse set of watersheds. Using physical models (such as global climate models or global hydrological models) coupled with LSTM [256] is proposed for long-term streamflow projections. The LSTM model is considered a post-processor that aims to constrain (i.e., reduce the uncertainty of) streamflow simulations from the physics-based model. To enhance streamflow prediction in the weather research and forecasting hydrological modeling system (WRF-Hydro), a method called WRF-Hydro-LSTM [257] is proposed. It utilizes LSTM to predict the residual errors of WRF-Hydro.
However, due to the presence of data errors in hydrological records, an unconstrained ML (such as LSTM) performs better than a PeML model (such as MC-LSTM) because of the ability of PDgML to learn and account for data errors [254]. In addition, three LSTM daily streamflow prediction models [258] are built and compared, including a LSTM post-processor trained on the United States National Water Model (NWM) outputs, a LSTM post-processor trained on the NWM outputs and atmospheric forcings, and a LSTM model trained only on atmospheric forcing. It concludes that a LSTM model trained only on atmospheric forcing outperforms the other two LSTMs in ungauged basins. Hybrid models may reduce the reliability of the model in some cases. Thus, PiML, PeML, and PaHL approaches for the reliability of rainfall-runoff forecasts require further development in the future.
Missing Data and Ungauged Basins. See Fig. 5 (label 1.5). Hydrologic predictions at many watersheds, such as rural watersheds, as well as cold and ungauged regions, are important but also challenging due to data scarcity [278]. Thus, many PaML-based methods are developed for poorly gauged or ungauged basins. For example, geo-spatio-temporal mesoscale data and attention-based deep learning [259] are used to improve long-term streamflow prediction. This study demonstrates that the 3D-CNN–Transformer achieves the best performance compared with TD-CNN–Transformer, TD-CNN–LSTNet, and 3D-CNN–LSTNet, and the application of hydroclimate data with proposed integrated networks is particularly useful for poorly gauged basins. In addition, while deep neural networks demonstrated stellar performance for streamflow predictions, performance nonetheless declined for prediction in ungauged regions. In order to reduce such errors, an input-selection ensemble method [260] is proposed by leveraging the deep network’s flexibility to integrate satellite-based soil moisture products or daily flow distributions. Furthermore, hybrid models can be suitable alternatives to hydrological models, particularly in watersheds where there is a lack of measured data (e.g., climatic parameters, land cover-plant growth data, soil data, stream conditions, and properties of groundwater aquifers), provided that appropriate inputs (such as rainfall) are used. Mohammadi et al. [196] show that AI-based hybrid models have generally led to more accurate streamflow estimates compared with conceptual rainfall-runoff models in ungauged basins. To enhance out-of-distribution prediction in data-scarce basins, some serial methods for developing the hybrid LSTM [261, 262], i.e., using the physics-based hydrologic model simulation outputs as another set of inputs in the LSTM network, are proposed. Additionally, to reproduce long-duration daily streamflow in poorly gauged watersheds, the transfer learning technique is employed in a pre-trained SWAT-LSTM model [233].
However, these PaML-based approaches are still only preliminary attempts. Deeply exploiting the transferability of ML models (such as domain adaptation methods [279], meta-learning methods [280], and pre-trained large models [foroumandichatgpt]) and the interpretability of physical models will effectively solve hydrological prediction problems in ungauged basins.
Parameterization. See Fig. 5 (label 1.6). Parameters for rainfall-runoff forecasts are not often easily measurable; rather they are estimated indirectly through either prior knowledge or model calibration. Calibration is used to estimate the model parameters by matching the model outputs at specific locations where the equivalent observation is available. Specifically, a ML approach and remotely sensed underlying surface data [263] are employed to establish models for estimating the runoff routing parameter. This study demonstrates that parameters estimated by ML from large samples have a robustness that can guarantee the overall performance of physical models. A deep neural network [264] is used to construct inverse mapping from informative responses to each of the selected parameters. Yang et al. [265] calibrate three key parameters of the Variable Infiltration Capacity model at every 1/8°$\degree$ grid-cell using ML-based maps (shuffled complex evolution algorithm). Furthermore, some PaHL-based methods, such as a differentiable programming framework [202] and differentiable parameter learning [243], are proposed to efficiently learn a mapping between inputs and parameters of process-based models in an end-to-end manner.
Parameterization is an important process of rainfall-runoff forecasts. Digging deeper into the physical properties of hydrological processes (such as mass conservation), we can achieve fast and effective parameterization based on the PiML, PeML, and PaHL technologies.
Uncertainty Estimation in Rainfall-runoff Forecasts. See Fig. 5 (label 1.7). Hydrological forecasts are often inaccurate because models suffer from inadequate conceptualization of underlying physics, inaccurate predictions of ML models, non-uniqueness of model parameters, and uncertainties in data used for calibration, (e.g., output data such as streamflow). Thus, uncertainty estimations are critical for actionable hydrological prediction. However, the majority of PaML-based rainfall-runoff studies do not provide uncertainty estimates. In order to address this problem, some PDgML approaches are proposed for model uncertainty, such as a data-driven approach [266], Bayesian processor [267], a hybrid ensemble and variational data assimilation framework [268], a hierarchical model tree [269], and probabilistic deep ensembles [270]. In addition, several LSTM-based approaches are presented for uncertainty estimation. Two strategies [271] are tested for using an LSTM in combination with Gaussian processes. In one strategy, the LSTM is used to parameterize a Gaussian process, and in the second, the LSTM is used as a forecast model with a Gaussian process post-processor. Althoff et al. [272] explore the predictive uncertainty in hydrological models based on LSTM by comparing a multiparameter ensemble to a dropout ensemble. Song et al. [274] propose a framework to quantify the uncertainty contributions of the sample set, ML approach, ML architecture, and their interactions to multi-step time-series forecasting based on the analysis of variance theory. Furthermore, to reduce the gap between ML and the hydrological sciences community for uncertainty estimations, Klotz et al. [273] establish an uncertainty estimation benchmarking procedure and present four PDgML (LSTM) baselines. Three baselines are based on mixture density networks, and one is based on Monte Carlo dropout. In addition, by integrating physics and ML approaches, PaHL has also been gaining attention, as it can alleviate the computational burden and improve efficiency for uncertainty estimation. For example, the LSTM model is considered a post-processor [256] that aims to reduce the uncertainty of streamflow simulations from the physics-based model. Serial hybrid LSTM [262] is introduced for uncertainty quantification. An integrated model (XAJ-MCQRNN) [275] that incorporates the Xinanjiang conceptual model and composite quantile regression NN is developed to overcome the phenomena of error propagation and accumulation encountered in multi-step-ahead flood probability density forecasts. In addition, a hybrid error correction model [276] is proposed for improving streamflow forecast accuracy. The hybrid model integrates the HBV model with the ML algorithm through a data assimilation technique.
In general, the integration of physics knowledge into ML for uncertainty estimation has the potential to enable better characterization of uncertainty. For example, ML surrogate models run the risk of producing physically inconsistent predictions; incorporating elements of physics could help with this issue. In addition, the reduced data needs of ML due to constraints for adherence to known rainfall-runoff processes could alleviate some of the high computational costs of NNs.
3.2 Hydrodynamic Process Understanding
Hydrodynamic models are mathematical models that attempt to replicate fluid motion and typically require solving computationally. These models simulate water motion by solving equations derived from the application of fundamental laws of physics, encompassing principles governing the conservation of mass, momentum, and energy. Hydrodynamic models can realize the simulation of hydrological processes in a time scale of seconds, hours, or days.
3.2.1 Physics in Hydrodynamic modeling
Depending on the spatial representation of the flow, hydrodynamic models can be dimensionally grouped into one-dimensional (1D) [281], two-dimensional (2D) [282], and three-dimensional (3D) models [283].
1D models. 1D hydrodynamic models have been widely used in modeling water flows [284]. Many hydraulic situations can make the 1D assumption, either because a more detailed solution is unnecessary (e.g., the purpose does not require knowledge in other dimensions) or because the flow is markedly 1D, such as in a confined channel or in a channel flow. Generally, 1D models solve equations derived by ensuring mass and momentum conservation, which yields the well-known 1D Saint-Venant equations,
$\begin{gathered}\frac{\partial h}{\partial t}+u\frac{\partial h}{\partial x}=q,\\ \frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+g\frac{\partial h}{\partial x}+g\left(S_{f}-S\right)=0,\end{gathered}$ (3)
where $x$ is the distance along the river channel; $t$ is time; $q$ is the water inflow per unit length of the channel from land surface and subsurface runoff, groundwater, and precipitation; $u$ and $h$ are the dynamics of velocity and water depth along the river channel, respectively; $g$ is gravity; $S$ is riverbed slope; and $S_{f}$ is the friction slope, which can be calculated using the Chezy–Manning equation [285].
2D models. The 2D models represent water or flood flow as a two-dimensional field with the assumption that the third dimension, water depth, is shallow in comparison to the other two dimensions [286, 287]. Most approaches solve two-dimensional shallow water equations, which represent mass and momentum conservation in a plane, and can be obtained by depth-averaging the Navier-Stokes equations,
$\begin{gathered}\frac{\partial h}{\partial t}+\frac{\partial q_{x}}{\partial x}+\frac{\partial q_{y}}{\partial y}=0\\ \underbrace{\frac{\partial q_{x}}{\partial t}}_{\begin{subarray}{c}\text{ local }\\ \text{ acceleration }\end{subarray}}+\underbrace{\frac{\partial}{\partial x}\left(uq_{x}\right)+\frac{\partial}{\partial y}\left(vq_{x}\right)}_{\begin{array}[]{c}\text{ convective }\\ \text{ acceleration }\end{array}}+\underbrace{gh\frac{\partial(h+z)}{\partial x}}_{\begin{array}[]{c}\text{ pressure }+\\ \text{ bed gradients }\end{array}}+\underbrace{\frac{gh^{2}\|\mathbf{q}\|q_{x}}{h^{7/3}}}_{\text{friction }}=0\\ \underbrace{\frac{\partial q_{y}}{\partial t}}_{\begin{array}[]{c}\text{ local}\\ \text{ acceleration }\end{array}}+\underbrace{\frac{\partial}{\partial y}\left(vq_{y}\right)+\frac{\partial}{\partial x}\left(uq_{y}\right)}_{\begin{array}[]{c}\text{ convective }\\ \text{ acceleration }\end{array}}+\underbrace{gh\frac{\partial(h+z)}{\partial y}}_{\begin{array}[]{c}\text{ pressure }+\\ \text{ bed gradients }\end{array}}+\underbrace{\frac{gn^{2}\|\mathbf{q}\|q_{y}}{h^{7/3}}}_{\text{friction }}=0,\end{gathered}$ (4)
where $x$ and $y$ are the two Cartesian directions; $q_{x}$ and $q_{y}$ are the $x$ and $y$ components of the discharge per unit width vector $\mathbf{q}$; $u$ and $v$ are the $x$ and $y$ components of the flow velocity, respectively; $z$ is the bed elevation; and $n$ is the Manning’s friction coefficient. In flood events, the key simplification introduced by the local inertial approximation is based on the assumption that the convective acceleration terms are negligible compared with the other terms and, therefore, can be ignored.
3D models. For many scales of water flow, complex 3D representation of flow dynamics has tended to be regarded as unnecessary, as a 2D shallow water approximation may be adequate [288]. However, modeling vertical turbulence, vortices, and floods occurring due to dam breaks is important. 3D models are therefore developed to allow the representation of vertical features. Some of these models solve the horizontal flow with 2D shallow water equations and include a quasi-3D extension to model velocity in vertical layers. Other 3D models are derived from the 3D Navier-Stokes equations, which describe the motion of fluid substances and are usually written as
$\begin{gathered}\frac{\partial u}{\partial t}+u\cdot\nabla u+\frac{1}{\rho}\nabla p=g+\mu\nabla\cdot\nabla u,\\ \text{ Incompressibility condition:}\quad\nabla\cdot u=0,\end{gathered}$ (5)
where $u$ is the velocity; $\rho$ is the fluid density; $p$ is pressure; $g$ is gravitational acceleration; and $\mu$ is kinematic viscosity. The incompressibility condition is based on the assumption that material density is constant within a fluid parcel. Depending on the representation of the processes, these models may be divided into two broad categories: Eulerian (grid-based) and Lagrangian (particle-based).
The following sections: Hydrodynamic Modeling Meets Physics-aware Machine Learning, Proposed HydroPML-based Flood Forecasting (Fig. 6), and Summary and Future Directions will be made available after the paper is accepted.
Acknowledgements
The work is jointly supported by the German Federal Ministry of Education and Research (BMBF) in the framework of the international future AI lab "AI4EO – Artificial Intelligence for Earth Observation: Reasoning, Uncertainties, Ethics and Beyond" (grant number: 01DD20001) and the project Inno__\__MAUS (grant number: 02WEE1632B), by German Federal Ministry for Economic Affairs and Climate Action in the framework of the "national center of excellence ML4Earth" (grant number: 50EE2201C), by the Excellence Strategy of the Federal Government and the Länder through the TUM Innovation Network EarthCare and by Munich Center for Machine Learning.